# A finite volume Simo-Reissner beam method for moored floating body dynamics

### v1

@@ 3.2 Numerical model
Wave gauge 2 0.0545 0.0589 0.0588 0.0638
Wave gauge 3 0.0576 0.0590 0.0626 0.0642
Wave gauge 4 0.0534 0.0589 0.0572 0.0636
Wave gauge 5 0.0539 0.0589 0.0571 0.0645
Wave gauge 6 0.0528 0.0597 0.0566 0.0642
Average diff. (%) 8.39 9.81
Max diff. (%) 13.07 13.45
The FFT spectrum is presented in Figure 14 for both H12T20 and H15T18 cases. These plots show that the dominant wave frequency observed in the experimental data is also predicted by the numerical model, indicating that the primary wave characteristics are well reproduced. Among all the wave gauge locations, only wave gauge 2 and 6 are shown here for brevity. These two locations were selected to highlight the performance of the model in regions with different mesh resolutions and wave conditions.
@@ 3.3 Floating Box Motions
Figure 15 compares the surge, heave, and pitch responses of the floating box for cases H12T20 and H15T18. It compares results from the present model against MoorDyn [9], and experimental data. A broad consistency exists between the numerical and experimental results for heave and surge motion. Although surge motion is overpredicted in both wave cases, the steeper wave condition (H15T18) shows a smaller phase shift relative to the experiments for both numerical models. The interpretation also applies to heave motions (Figures 15(c) and 15(d)).
@@
Amplitude is computed as the average of multiple peak-to-trough values in the time interval between 8 and 16 seconds, corresponding to approximately 4.4 wave periods for case H15T18 and 4 wave periods for case H12T20. This provides a consistent estimate of the characteristic oscillation magnitude. Percentage error is reported relative to the experimental amplitude. For H12T20, both the present model and MoorDyn overpredict the surge amplitude (Figure 15(a)), with the present model deviating by 28.69% and MoorDyn by 27.40%. In heave, both models are within a close range of the experimental results (Figure 15(c)), with the present model showing a deviation of 6.46% and MoorDyn 7.03%. For pitch, both numerical models slightly underestimate the amplitude; MoorDyn is slightly closer to the experimental value (Figure 15(e)), with a 3.93% error compared to 5.23% from the present model. In the steeper wave case H15T18, the present model performs slightly better in heave and pitch predictions. In heave, the present model has a deviation of 6.19%, compared to 7.76% for MoorDyn (Figure 15(d)). For pitch, the present model shows 4.10% error, while MoorDyn underpredicts the amplitude with a larger deviation of 38.60% (Figure 15(f)). These findings indicate that the present model captures pitch responses more consistently under steeper wave conditions (H15T18) than MoorDyn. Stronger nonlinear coupling and closer alignment between the wave period and the system’s natural pitch frequency likely contribute to the lower pitch prediction error in H15T18 compared to H12T20. Nevertheless, further investigation would help confirm and refine these observations.
This discrepancy in pitch motion is also evident in Figures 15(e) and 15(f). One possible factor causing this offset is the wooden plate attached to the box’s front face and the reflective markers attached. This wooden plate is not present in numerical simulation, which can lead to some discrepancies in pitch motion. Moreover, a more important note from the experimental report is that during the experiments, there was splashing on this plate [22, 50], which may explain the differences in pitch motion compared to measured data. Additionally, previous numerical studies [9, 22] have also reported similar offsets in pitch motion.
@@
Experimental MoorDyn [9] Present Model
@@
Surge (m) 0.22 - 0.28 27.40 0.29 28.69
Heave (m) 0.11 - 0.12 7.03 0.12 6.46
Pitch (°) 9.0 - 8.71 3.93 8.59 5.23
@@
Surge (m) 0.23 - 0.26 11.76 0.27 15.36
Heave (m) 0.14 - 0.13 7.76 0.13 6.19
Pitch (°) 13.40 - 8.23 38.60 13.95 4.10
@@ 3.4 Mooring Line tensions
The time series of mooring line responses for the two wave cases are shown in Figure 16. Since the system is a four-point symmetric mooring system, only the tensions for line 1 and line 3 at the anchors are shown (see Figure 7). The simulations and experiments show a close agreement in the shape of the tension time history. It is evident that the waveward cables (line 1) experience larger tension offsets compared to the leeward cables (line 3) in both cases. This is expected, as waveward cables are more influenced by mean wave drift forces, leading to occasional snap loads.
A quantitative comparison of mooring line tension amplitudes between the present model, MoorDyn, and experimental results is provided in Table 8. Tension amplitude is computed as the average of multiple peak-to-trough values in the time interval between 8 and 16 seconds, covering approximately 4.4 wave periods for case H15T18 and 4 wave periods for case H12T20. This approach provides a reliable estimate of the dominant dynamic response. Percentage error is reported relative to the experimental amplitude. For H12T20, the present model slightly overpredicts line 1 tension (Figure 16(a)) by 6.87% compared to the experimental value. MoorDyn, on the other hand, exhibits a significantly higher deviation of 28.17%. In line 3 (Figure 16(c)), both the present model and MoorDyn show reasonable agreement with experiments, though MoorDyn tends to underestimate the amplitude by 13.62% as compared to the present model which shows a deviation of 3.17%. In the steeper wave case (H15T18), the discrepancies become more pronounced. The present model underestimates the line 1 amplitude (Figure 16(b)) by 32.68%, while MoorDyn shows a slightly larger deviation of 36.45%. In line 3 (Figure 16(d)), the present model exhibits a 23.12% error, whereas MoorDyn deviates by 39.68%, again showing a noticeably larger mismatch with the experimental data. These results suggest that the present model may offer an improvement across different wave conditions. This mainly could be related to the Simo-Reissner beam formulation providing a more detailed representation of the internal forces under dynamic deformation, which may contribute to slightly closer agreement with the experimental measurements in mooring tensions.
@@
MoorDyn [9] 1.7481 28.17 1.3121 13.62
Present Model 1.4576 6.87 1.4709 3.17
@@
Experimental 2.4585 - 1.8764 -
MoorDyn [9] 1.5623 36.4531 1.1317 39.6876
Present Model 1.655 32.6825 1.4425 23.1240
Several factors may contribute to the discrepancies observed in the predicted mooring line tensions. While mean wave drift forces likely account for part of the offset, additional modelling assumptions could also influence the results. These include the simplified representation of the mooring line connections, particularly the absence of flexibility or deformation at the chain links, which were present in the experimental setup but are not captured in the current model. Assumed values for axial stiffness and hydrodynamic coefficients, such as drag and added mass, introduce further uncertainty. Moreover, the assumption of a quiescent fluid surrounding the cables neglects the influence of local flow disturbances caused by the floating body and incoming waves. This may be particularly important in the steeper wave case (H15T18), where more energetic interactions occur near the fairlead. Similar limitations have been discussed in prior numerical studies [9, 22]. While these explanations are plausible, isolating their effects remains difficult without dedicated experiments or higher-fidelity coupling. This underscores the inherent challenges of reproducing detailed experimental conditions in numerical simulations.
@@ 4 Conclusion
This paper presents a novel unified finite-volume-based discretization framework for modeling mooring cable dynamics, integrating multiphase CFD flow simulations and a six-degree-of-freedom motion solver. The framework combines OpenFOAM’s native six-degree-of-freedom motion solver with a finite-volume-based Simo-Reissner beam solver. This integration enables detailed analysis of the moored motion of a floating rigid box, accounting for the bending, shear, and torsional deformations of the mooring line. To validate the proposed approach, numerical results of the current model for a floating rigid box moored with four catenary lines are compared to both experimental and numerical data to assess the framework’s ability to capture complex interactions between the structure and its environment. The following observations are made from the numerical analyses:
@@
  • Aligning the first wave peak with the experimental data removed an initial phase lag in the surface elevation time histories and simplified direct comparisons, demonstrating that the numerical model successfully reproduces the dominant wave frequency and closely matches the amplitudes reported in Wu et al. [50] (see Figure 14). However, the coarser mesh region near the outlet exhibits a noticeable offset in wave amplitude (Figures 13(e) and 13(f)). This confirms the model’s ability to resolve local wave behaviour near the structure but highlights limitations in far-field accuracy.
@@
  • Mooring line tensions show a strong correlation with experimental data on the leeward side, with both numerical models underpredicting the waveward cable tensions. However, the present model consistently provides a closer match to experimental values compared to MoorDyn, as seen in line 1 of H12T20 (6.87% vs 28.17%), line 3 of H12T20 (3.17% vs 13.62%), and line 3 of H15T18 (23.12% vs 39.68%), where MoorDyn exhibits substantially larger deviations. The improved accuracy of the present model suggests a better representation of mooring line dynamics under wave-induced forces. The remaining discrepancies are primarily attributed to the influence of mean wave drift forces, which are not fully captured in the numerical models, as well as additional assumptions regarding axial stiffness, fairlead and anchor connections, and the quiescent fluid assumption. These factors collectively contribute to the observed offset. These results indicate that the present model enhances the reliability of mooring load predictions.
@@
  • Surge and heave motions show good agreement with experimental data, with both the present model and MoorDyn slightly overpredicting these motions. In surge, MoorDyn exhibits a smaller deviation than the present model, for both H12T20 and H15T18 cases. In heave, both models closely match the experimental amplitude for H12T20, with the present model deviating by 6.46% and MoorDyn by 7.03%. For H15T18, both models slightly underpredict the motion, with the present model showing a smaller error of 6.19% compared to 7.76% for MoorDyn. (see Figures 15(a), 15(c), 15(c), 15(d)). This supports the validity of the present model in predicting primary translational motions of the floating body.
@@
  • The largest discrepancies arise in pitch motion (see Figures 15(e) and 15(f)), where the present model matches the observed amplitude more consistently under the steeper wave condition (H15T18). This improved performance may be linked to the closer alignment between the wave period and the system’s natural pitch frequency, which tends to amplify any modeling inaccuracies or strengths. By contrast, MoorDyn underpredicts the amplitude in these conditions. Some of the remaining errors appear to stem from experimental uncertainties, such as a wooden plate on the box’s front face, which is omitted in the numerical model. Similar discrepancies have also been noted in prior studies [9, 22], highlighting ongoing challenges in numerically simulating moored floating structures. This indicates that the present model better captures rotational dynamics under nonlinear wave-structure interaction.
@@
  • The proposed model demonstrates efficient scaling (see Figure LABEL:fig:strongScaling), with a notable decrease in execution time as additional CPU cores are utilised. However, beyond 256 CPU cores, the improvement in performance becomes less significant. This suggests the model is suitable for large-scale simulations but benefits taper with excessive parallelisation.
@@
Declaration of generative AI and AI-assisted technologies in the writing process During the preparation of this work the author(s) used Grammarly and ChatGPT to improve the grammar and readability of the manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.

### v2

@@ 3.2 Numerical model
Wave gauge 2 0.0546 0.0541 0.0604 0.0601
Wave gauge 3 0.0539 0.0535 0.0605 0.0599
Wave gauge 4 0.0534 0.0533 0.0591 0.0588
Wave gauge 5 0.0533 0.0534 0.0593 0.0590
Wave gauge 6 0.0528 0.0540 0.0576 0.0599
Average diff. (%) 0.86 1.30
Max diff. (%) 2.27 4.00
The FFT spectrum is presented in Figure 14 for both H12T20 and H15T18 cases. These plots show that the dominant wave frequency observed in the experimental data is also predicted by the numerical model, indicating that the primary wave characteristics are well reproduced. Among all the wave gauge locations, only wave gauges 2 and 6 are shown here for brevity. These two locations were selected to highlight the performance of the model in regions with different mesh resolutions and wave conditions.
@@ 3.3 Floating Box Motions
Figure 15 compares the surge, heave, and pitch responses of the floating box for cases H12T20 and H15T18. It compares results from the present model against MoorDyn [chen_cfd_2022], and experimental data. The numerical time series were shifted in time so that their first peak coincides with the experimental record, thereby removing the initial phase lag and making the comparison clearer. A broad consistency exists between the numerical and experimental results for heave and surge motion. Although surge motion is overpredicted in both wave cases, the steeper wave condition (H15T18) shows a smaller phase shift relative to the experiments for both numerical models. The interpretation also applies to heave motions (Figures 15(c) and 15(d)).
@@
Amplitude is computed as the average of multiple peak-to-trough values in the time interval between 8 and 16 seconds, corresponding to approximately 4.4 wave periods for case H15T18 and four wave periods for case H12T20. This provides a consistent estimate of the characteristic oscillation magnitude. The percentage error is reported relative to the experimental amplitude. For H12T20, both the present model and MoorDyn overpredict the surge amplitude (Figure 15(a)), with the present model deviating by 9.09% and MoorDyn by 27.40%. In heave, the present model matches the experimental data more closely than MoorDyn (Figure 15(c)), with a deviation of 3%, compared with 9.08% for MoorDyn. For pitch, both numerical models slightly underestimate the amplitude; the present model is closer to the experimental value (Figure 15(e)), with a 2.55% error compared to 3.93% from the MoorDyn. In the steeper wave case (H15T18), the present model performs better in all surge, heave and pitch predictions. For surge, the present model differs from the experiments by 8.6%, while MoorDyn shows a larger 13.04% error. In heave, the present model has a deviation of 6.49%, compared to 7.14% for MoorDyn (Figure 15(d)). For pitch, the present model shows 4.32% error, while MoorDyn underpredicts the amplitude with a larger deviation of 38.60% (Figure 15(f)). These findings indicate that the present model captures pitch responses more consistently under steeper wave conditions (H15T18) than MoorDyn. Stronger nonlinear coupling and closer alignment between the wave period and the system’s natural pitch frequency likely contribute to the lower pitch prediction error in H15T18 compared to H12T20. Overall, the present model delivers smaller amplitude errors than MoorDyn across both wave cases, demonstrating improved predictive capability. Nevertheless, further investigation would help confirm and refine these observations.
Discrepancies between the numerical and experimental motion responses, particularly in pitch and surge, can be attributed to several factors. One likely contributor is the wooden plate with reflective markers mounted on the front face of the floating box in the experimental setup [wu_experimental_2019], which was not included in the numerical model. This addition introduces asymmetry that could influence the added mass and hydrodynamic loading, especially in pitch. Moreover, splashing observed on this plate during the experiments [dominguez_sph_2019, wu_experimental_2019] may further affect the pitch response and tracking accuracy. In the present work, the plate has been explicitly accounted for in the reported mass properties (centre of gravity and moments of inertia) to represent its static contribution, but no additional empirical damping was introduced, since such effects cannot be quantified without dedicated decay tests. This choice preserves the physical transparency of the numerical model, avoiding case specific calibration. The numerical model also assumes ideal second-order Stokes waves, while in the physical setup, the wave maker introduces startup transients and small irregularities, potentially contributing to phase shifts or amplitude deviations. Additional simplifications, such as the assumption of quiescent flow around the mooring lines and approximated hydrodynamic coefficients, may also influence the simulated responses. Similar discrepancies have been reported in previous numerical studies using this benchmark case [chen_cfd_2022, dominguez_sph_2019].
@@
Experimental MoorDyn [chen_cfd_2022] Present Model
@@
Surge (m) 0.22 - 0.28 27.40 0.24 9.09
Heave (m) 0.11 - 0.12 9.08 0.1133 3
Pitch (°) 9.0 - 8.71 3.93 8.77 2.55
@@
Surge (m) 0.23 - 0.26 13.04 0.25 8.6
Heave (m) 0.14 - 0.13 7.14 0.131 6.49
Pitch (°) 13.40 - 8.23 38.60 13.98 4.32
@@ 3.4 Mooring Line tensions
The time series of mooring line responses for the two wave cases are shown in Figure 16. Since the system is a four-point symmetric mooring system, only the tensions for line 1 and line 3 at the anchors are shown (see Figure 7). The simulations and experiments show a close agreement in the shape of the tension time history. The waveward cables (line 1) experience larger tension offsets compared to the leeward cables (line 3) in both cases. This is expected, as waveward cables are more influenced by mean wave drift forces, leading to occasional snap loads.
A quantitative comparison of mooring line tension amplitudes between the present model, MoorDyn, and experimental results is provided in Table 8. Tension amplitude is computed as the average of multiple peak-to-trough values in the time interval between 8 and 16 seconds, covering approximately 4.4 wave periods for case H15T18 and four wave periods for case H12T20. This approach provides a reliable estimate of the dominant dynamic response. The percentage error is reported relative to the experimental amplitude. For H12T20, the present model slightly underpredicts line 1 tension (Figure 17(a)) by 3.18% compared to the experimental value. MoorDyn, on the other hand, exhibits a significantly higher deviation of 28.17%. In line 3 (Figure 16(c)), both the present model and MoorDyn show reasonable agreement with experiments, though MoorDyn tends to underestimate the amplitude by 13.62% as compared to the present model, which shows a deviation of 7.43%. In the steeper wave case (H15T18), the discrepancies become more pronounced. The present model underestimates the line 1 amplitude (Figure 17(b)) by 23.62%, while MoorDyn shows a slightly larger deviation of 26.94%. In line 3 (Figure 16(d)), the present model exhibits a 25.25% error, whereas MoorDyn deviates by 39.68%, again showing a noticeably larger mismatch with the experimental data. These results suggest that the present model may offer an improvement across different wave conditions. This improvement may be mainly due to the Simo-Reissner beam formulation providing a more detailed representation of the internal forces under dynamic deformation, which may contribute to slightly closer agreement with the experimental measurements in mooring tensions. It is worth noting that no snap load events were detected in the tested cases. The selected catenary configuration and the range of wave heights (0.12–0.15 m) ensured that the mooring lines remained continuously under tension throughout the simulations.
@@
MoorDyn [chen_cfd_2022] 1.7481 28.17 1.3121 13.62
Present Model 1.3205 3.18 1.406 7.43
@@
Experimental 2.1385 - 1.8764 -
MoorDyn [chen_cfd_2022] 1.5623 26.94 1.1317 39.6876
Present Model 1.633 23.624 1.4025 25.25
Several factors may contribute to the discrepancies observed in the predicted mooring line tensions. While mean wave drift forces likely account for part of the offset, additional modelling assumptions could also influence the results. These include the simplified representation of the mooring line connections, particularly the absence of flexibility or deformation at the chain links, which were present in the experimental setup but are not captured in the current model. Assumed values for axial stiffness and hydrodynamic coefficients, such as drag and added mass, introduce further uncertainty. Moreover, the assumption of a quiescent fluid surrounding the cables neglects the influence of local flow disturbances caused by the floating body and incoming waves. This may be particularly important in the steeper wave case (H15T18), where more energetic interactions occur near the fairlead. Similar limitations have been discussed in prior numerical studies [chen_cfd_2022, dominguez_sph_2019]. While these explanations are plausible, isolating their effects remains difficult without dedicated experiments or higher-fidelity coupling. This underscores the inherent challenges of reproducing detailed experimental conditions in numerical simulations.
@@
To investigate the sensitivity of mooring line tension to axial stiffness, the parameter $EA$ was varied from the baseline value of $19$ N to $9.5$ N and $38$ N for both H12T20 and H15T18 cases. The comparison in Figure 17 shows that in both wave conditions, reducing the stiffness to $9.5$ N results in a visibly damped response with lower peak tensions and broader force profiles, while increasing it to $38$ N produces sharper and more pronounced peaks. The baseline value of $EA=19$ N consistently provides the best agreement with experimental results across both cases, especially in terms of peak magnitude and phase alignment. The H15T18 case, featuring a steeper wave profile and more pronounced wave loads, demonstrates greater sensitivity to $EA$ changes. Here, as also reported by chen_cfd_2022, increasing $EA$ does help in partially recovering the under-predicted tension peaks. However, the current results show that overly stiff lines ($EA=38$ N) tend to overshoot the experimental tension, introducing unrealistic spikes. These results show that axial stiffness is really a property of the whole mooring system, not just the cable material. Chain-link flexibility, contact with the seabed or fairlead, and hydrodynamic damping all change the stiffness that the line presents in practice. Calibrating $EA$ against integrated dynamic responses rather than static material tests is therefore crucial for reliable modelling of mooring line behaviour.
(a) line1 Tension - H12T20
(b) line1 Tension - H15T18
Figure 17: Influence of axial stiffness(EAEA) on mooring line tension
@@ 4 Conclusion
This paper presents a novel unified finite-volume-based discretisation framework for modelling mooring cable dynamics, integrating multiphase CFD flow simulations and a six-degree-of-freedom motion solver. The framework combines OpenFOAM’s native six-degree-of-freedom motion solver with a finite-volume-based Simo-Reissner beam solver. This integration enables detailed analysis of the moored motion of a floating rigid box, accounting for the bending, shear, and torsional deformations of the mooring line. To validate the proposed approach, numerical results of the current model for a floating rigid box moored with four catenary lines are compared to both experimental and numerical data to assess the framework’s ability to capture complex interactions between the structure and its environment. The following observations are made from the numerical analyses:
@@
  • Aligning the first wave peak with the experimental data removed an initial phase lag in the surface elevation time histories and simplified direct comparisons, demonstrating that the numerical model successfully reproduces the dominant wave frequency and closely matches the amplitudes reported in wu_experimental_2019 (see Figure 14). However, the coarser mesh region near the outlet exhibits a noticeable offset in wave amplitude (Figures 13(e) and 13(f)). This confirms the model’s ability to resolve local wave behaviour near the structure but highlights limitations in far-field accuracy.
@@
  • Mooring line tensions on the leeward side show a strong correlation with the experimental data, while both numerical models deviate from the experimental values on the waveward side. However, the present model consistently provides a closer match to experimental values compared to MoorDyn, as seen in line 1 of H12T20 (3.18% vs 28.17%), line 3 of H12T20 (7.43% vs 13.62%), and line 3 of H15T18 (25.25% vs 39.68%), where MoorDyn exhibits substantially larger deviations. The improved accuracy of the present model suggests a better representation of mooring line dynamics under wave-induced forces. The remaining discrepancies are primarily attributed to the influence of mean wave drift forces, which are not fully captured in the numerical models, as well as additional assumptions regarding axial stiffness, fairlead and anchor connections, and the quiescent fluid assumption. These factors collectively contribute to the observed offset. These results indicate that the present model enhances the reliability of mooring load predictions.
@@
  • Surge and heave motions show good agreement with experimental data, with both the present model and MoorDyn slightly overpredicting these motions. In surge, the present model shows smaller deviations than MoorDyn in both cases, with 9.09 % compared to 27.40% for H12T20 and 8.6% compared to 13.04% for H15T18. In heave, both models closely match the experimental amplitude for H12T20, with the present model deviating by 3% and MoorDyn by 9.08%. For H15T18, both models slightly underpredict the heave motion, with the present model showing a smaller error of 6.49% compared to 7.14% for MoorDyn. (see Figures 15(a), 15(c), 15(c), 15(d)). This supports the validity of the present model in predicting primary translational motions of the floating body.
@@
  • The largest discrepancies arise in pitch motion (see Figures 15(e) and 15(f)), where the present model matches the observed amplitude more consistently under the steeper wave condition (H15T18). This improved performance may be linked to the closer alignment between the wave period and the system’s natural pitch frequency, which tends to amplify any modelling inaccuracies or strengths. By contrast, MoorDyn underpredicts the amplitude in these conditions. Some of the remaining errors appear to stem from experimental uncertainties, such as a wooden plate on the box’s front face, which is omitted in the numerical model. Similar discrepancies have also been noted in prior studies [chen_cfd_2022, dominguez_sph_2019], highlighting ongoing challenges in numerically simulating moored floating structures. This indicates that the present model better captures rotational dynamics under nonlinear wave-structure interaction.
@@
  • The proposed model demonstrates efficient scaling (see Figure LABEL:fig:strongScaling), with a notable decrease in execution time as additional CPU cores are utilised. However, beyond 256 CPU cores, the improvement in performance becomes less significant. This suggests the model is suitable for large-scale simulations, but benefits wane as the number of cells per CPU core becomes small.
@@
Declaration of generative AI and AI-assisted technologies in the writing process During the preparation of this work, the author(s) used Grammarly and ChatGPT to improve the grammar and readability of the manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Appendix A Grid convergence study
Three different mesh resolutions were selected for the grid convergence study. For this study, the H12T20 case with nominal wave height ($H=12cm$) has been used. The medium-resolution mesh has a cell size of $0.01\times 0.011\times 0.01$ in the refined region of the domain. Outside this region, the mesh is gradually stretched using a grading ratio of $1.07$ (see Figure 8). The fine mesh has smaller cells than the medium mesh by a factor of $\sqrt{2}$, while the coarse mesh has larger cells by the same factor.
(a) Surge
(b) Heave
(c) Pitch
Figure 18: Comparing box motion using different grid resolutions
The motion of the box for three different grid sizes is illustrated in Figure 18. The surge motion, which is directly influenced by the restraining force of the mooring lines, remains nearly identical across all mesh resolutions. The results for heave and pitch differ across the coarse, medium, and fine meshes. In both cases, the medium-resolution results fall between those of the coarse and fine meshes, and are noticeably closer to the fine mesh results. This indicates that while there is some mesh sensitivity, the medium mesh provides a reasonable approximation of the converged solution.
Appendix B Time step sensitivity study
In addition to the grid convergence analysis presented in Appendix A, a time step sensitivity analysis was carried out to assess the influence of the Courant number on the accuracy of the floating body responses. Three cases were considered, corresponding to maximum Courant numbers of 0.25, 0.50, and 0.75. Since the solver employs an adaptive time step based on the Courant condition, these values directly control the maximum allowable time step size. The body motions obtained from these simulations were compared against the experimental measurements for the H12T20 case.
Figures 19(a)19(c) present the pitch, heave, and surge responses for the three Courant number values. Overall, the results indicate that the simulated motions are only marginally affected by the time step size. For pitch (Figure 19(a)), all three simulations follow the experimental trend with good phase alignment and comparable amplitudes, though the Courant = 0.25 case shows slightly sharper oscillations. For heave (Figure 19(b)), the three simulations are almost indistinguishable, all showing excellent agreement with the experimental data. Surge motion (Figure 19(c)) is similarly robust, with the Courant=0.25 and Courant=0.5 cases providing nearly identical predictions and the Courant=0.75 case showing only minor differences.
These results suggest that the present model is not particularly sensitive to the chosen Courant number within the tested range. Based on this analysis, a maximum Courant number of 0.50 was selected for the main simulations, as it provides a good balance between numerical accuracy and computational efficiency.
(a) Pitch
(b) Heave
(c) Surge
Figure 19: Time step sensitivity analysis (H12T20): simulated and experimental responses in pitch, heave, and surge.